The demand function of a commodity at a sales price p is given by: D(p) = 700/p − 250 p^2−0.002p Find the elasticity of demand and the price of unit elasticity(meaning what is p when |E(p)| = 1).


1
Expert's answer
2021-03-16T08:29:23-0400

"\frac{dD(p)}{dp}= -700p^{-2}-500p-0.002\\
\text{which is the elasticity of demand}\\
\text{The price of unit}.\\
E(p)= -700p^{-2}-500p-0.002\\
E(p)=1, \\
1=-700p^{-2}-500p-0.002\\
\text{multiply both sides by} p^2 \text{and collect like terms}\\
500p^3+1.002p^2+700=0\\"